OpenStudy (kj4uts):

f(x)=2x+1 and g(x)=x^2-7 find (fog)(2) Please explain how I solve a question like this, thank you!

OpenStudy (anonymous):

fog means really the composition of f with g (o is the composition operator). All that means is you evaluate g(2) first, and substitute that answer into f: g(2) = 2^2 - 7 =-3 f(-3) = 2(-3)+1 = -5 so fog(2) = -5